# On the Hurwitz Stability of Hurwitz-Type Matrix Polynomials

### v1

@@ 4 Bezoutian and HTM polynomials
$=$ ${\bf h}_{2m}^{*}(\bar{x}^{2})\left(x^{-2}I_{q},x^{-4}I_{q},\ldots\right)H^{(1)}_{\infty}\left(y^{-2}I_{q},y^{-4}I_{q},\ldots\right)^{\mathsf{T}}{\bf h}_{2m}(y^{2}).$
@@
${\bf h}_{2m}^{(11)}(\xi)+{\bf h}_{2m}^{(12)}(\xi)=$ $\left(\xi^{-2}I_{q},\xi^{-4}I_{q},\ldots\right)^{\mathsf{T}}{\bf h}_{2m}(\xi^{2}),$ (99)
${\bf h}_{2m}^{(21)}(\xi)+{\bf h}_{2m}^{(22)}(\xi)=$ $\left(\xi^{-1}I_{q},\xi^{-3}I_{q},\ldots\right)^{\mathsf{T}}{\bf h}_{2m}(\xi^{2}),$ (100)
${\bf g}_{2m+1}^{(11)}(\xi)+{\bf g}_{2m+1}^{(12)}(\xi)=$ $\left(I_{q},\xi^{-2}I_{q},\ldots\right)^{\mathsf{T}}{\bf g}_{2m+1}(\xi^{2}),$ (101)
${\bf g}_{2m+1}^{(21)}(\xi)+{\bf g}_{2m+1}^{(22)}(\xi)=$ $\left(\xi^{-1}I_{q},\xi^{-3}I_{q},\ldots\right)^{\mathsf{T}}{\bf g}_{2m+1}(\xi^{2}),$ (102)
@@
$\widehat{\bf h}_{2m}^{(11)},\,\widehat{\bf h}_{2m}^{(21)},\,\widehat{\bf g}_{2m+1}^{(11)}\quad\mbox{and}\quad\widehat{\bf g}_{2m+1}^{(11)}$ (108)
@@
$\widehat{\bf h}_{2m}^{(11)}(y)=\left(y^{-2}{\bf h}_{2m}(y^{2}),y^{-4}{\bf h}_{2m}(y^{2}),\ldots,y^{-2m}{\bf h}_{2m}(y^{2})\right)^{\mathsf{T}}$ (109)
@@
${\bf h}_{2m}^{(11)}(y)=\left(y^{-2}{\bf h}_{2m}(y^{2}),y^{-4}{\bf h}_{2m}(y^{2}),\ldots,y^{-2m}{\bf h}_{2m}(y^{2}),0_{q},0_{q}\ldots,\right)^{\mathsf{T}}.$
@@
is called a symmetrizer of $Q(x)$. In [51, page 455], the concept of symmetrizer is introduced for scalar polynomials. Additionally, let
@@
Let Condition C, given by (14), be satisfied. Furthermore, let ${\mathscr{G}}_{2m}^{(k)}$, ${\mathscr{G}}_{2m+1}^{(k)}$ for $k=1,2$ be as defined in (88) and (89). Additionally, let ${\bf h}_{n}$, ${\bf g}_{n}$, $S$, $\widetilde{J}_{m}$, $H_{1,j}$, $H_{2,j}$, $F_{0}^{2k}(x)$ and $F_{1}^{2k-1}(x)$ be as in (5), (9), (110), (103), (18), (111) and (112), respectively.
@@
${\mathscr{G}}_{2m}^{(2)}(x,y)=$ $-F_{1}^{(2m-1)^{*}}(\bar{x})S({\bf h}_{2m}^{*})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}S({\bf h}_{2m})F_{1}^{(2m-1)}(y).$ (114)
@@ 5 Hurwitness of Hurwitz-type matrix polynomials
In this section, we verify that an HTM polynomial ${\bf f}_{n}$, defined in Definition 1.1 and satisfying Condition C given by (14), is indeed a Hurwitz matrix polynomial. In [67], the Hurwitzness of HTM polynomials was considered. Unfortunately, the proof of [67, Theorem 3.5] regarding the inertia of the matrix polynomial ${\bf f}_{n}$ for $n=2m$ is incorrect, and the proof for $n=2m+1$ is absent. See Remark 5.8. Recall that the inertia of matrix polynomials, denoted by $\gamma({\bf f}_{n})$, provides the number of eigenvalues of ${\bf f}_{n}$ with positive, negative and real zero parts, respectively; see [53, page 410]. We rewrite Corollary 4.5 for the pair $(ix,iy)$ instead of $(x,y)$.
@@
$V_{[0,2m-1]}(x):=$ $\left(I_{q},\bar{x}I_{q},\bar{x}^{2}I_{q},\ldots,\bar{x}^{2m-2}I_{q},\bar{x}^{2m-1}I_{q}\right)^{\mathsf{*}},$ (130)
$V_{[0,2m]}(x):=$ $\left(I_{q},\bar{x}I_{q},\bar{x}^{2}I_{q},\ldots,\bar{x}^{2m-1}I_{q},\bar{x}^{2m}I_{q}\right)^{\mathsf{*}}.$ (131)
Furthermore, let
@@
$\begin{pmatrix}F_{0}^{(2m)}(x)\\ F_{1}^{(2m-1)}(x)\end{pmatrix}={\mathscr{J}}_{m-1}^{(m)}{\mathscr{T}}_{2m+1}V_{[0,2m]}(x),$ (133)
where ${\mathscr{T}}_{2m}$ (resp. ${\mathscr{T}}_{2m+1}$) are elementary block matrices. The following remark readily follows.
@@
Equalities (128) and (129) can be written as follows:
@@
Each HTM polynomial satisfying Condition C (14) is a Hurwitz matrix polynomial.
@@
A $q\times q$ HTM polynomial ${\bf f}_{2}$ of degree $2$ is a Hurwitz matrix polynomial, since ${\bf f}_{2}$ admits the presentation
${\bf f}_{2}(z)=I_{q}z^{2}+s_{0}z+s_{0}^{-1}s_{1}$
and it readily satisfies Condition C (14).
@@
a) the matrix $B_{x{\bf g}_{n}^{*},{\bf h}_{n}^{*}}(x{\bf g}_{n},{\bf h}_{n})$ is a positive definite, and
b) the matrix $B_{{\bf g}_{n}^{*},{\bf h}_{n}^{*}}({\bf g}_{n},{\bf h}_{n})$ is a negative definite.
@@
From Theorem 3.2 (b), see that assertion a) and b) are valid. ∎
@@
In [51, Theorem 2], it is proven that a scalar polynomial $f_{n}$ is Hurwitz if and only if $B_{xg_{n}^{*},h_{n}^{*}}(xg_{n},h_{n})$ is negative definite and $B_{g_{n}^{*},h_{n}^{*}}(g_{n},h_{n})$ is positive definite. Note that the Markov parameters related to [51, Theorem 2] have the opposite sign compared with those associated with parts (a) and (b) of Proposition 5.6.
Remark 5.8.
Let $L_{1}$ and $L$ be as in (136). In the proof of [67, Theorem 3.5] regarding the inertia representation of matrix polynomials ${\bf f}_{n}$, the [67, Equality (3.6)]
$B_{L_{1}^{*},L^{*}}(L,L_{1})=2iB_{{\bf h}_{2m}^{*}(-\bar{x}^{2}),x{\bf g}_{2m}^{*}(-\bar{x}^{2})}({\bf h}_{2m}(-x^{2}),-x{\bf g}_{2m}(-x^{2}))$ (139)
is employed for $n=2m$. Furthermore, from (137) and (124), we see that (139) is not correct. The proof of [67, Theorem 3.5] for $n=2m+1$ is not given.
In Theorem 4.10 of [67], a “stability criterion via Markov parameters” for matrix polynomials ${\bf f}_{n}$ in even and odd cases is suggested. We have the following remarks regarding this.
Remark 5.9.
a) The family of matrix polynomials considered in [67, Theorem 4.10] is related to the HTM polynomials. Other matrix polynomials, for example, the Hurwitz polynomial (143), are not covered by [67, Theorem 4.10].
b) The proof of [67, Theorem 4.10] uses a decomposition of a certain Bezoutian form as a polynomial in $z$ and $u$, without verifying whether Equality (73) is satisfied.
c) The proof of [67, Theorem 4.10] separately analyzes decompositions of the Bezoutian forms related to the odd part $F_{o}$ of the matrix polynomial $F$, and the even part $F_{e}$ of $F$, without indicating the relationship between the Bezoutian form of $F$ and those of $F_{o}$ and $F_{e}$.
d) The proof of [67, Theorem 4.10] is performed out for the case $n=2m+1$. The authors of [67] state that the case $n=2m$ reduces to [67, Theorem 4.4]. In turn, the proof of [67, Theorem 4.4] relies heavily on [67, Theorem 3.5]. From Remark 5.8, we know that the proof of [67, Theorem 3.5] is incorrect.
The following example demonstrates that there exist HTM polynomials that, despite not satisfying Condition C (14), are nevertheless Hurwitz polynomials.
Example 5.10.
Let the Markov parameters, also called moments, be given by
$s_{0}=\begin{pmatrix}4&i\\ -i&4\end{pmatrix},\quad s_{1}=\begin{pmatrix}1&i\\ -i&2\end{pmatrix},\quad s_{2}=\begin{pmatrix}3&i\\ -i&4\end{pmatrix}.$
The corresponding block Hankel matrix $H_{1,1}$ and $H_{2,0}$ are positive definite matrices. According to parts (b) and (f) of Theorem 3.2, the matrix
${\bf f}_{3}(z)=$ $(-1)(Q_{2,1}^{*}(-\bar{z}^{2})+z\,P_{2,1}^{*}(-\bar{z}^{2}))$
$=$ $Iz^{3}+A_{1}z^{2}+A_{2}z+A_{3}$
is an HTM polynomial, where
$A_{1}=\left(\begin{array}[]{cc}4&i\\ -i&4\\ \end{array}\right),A_{2}=\left(\begin{array}[]{cc}5&-2i\\ 2i&3\\ \end{array}\right),A_{3}=\left(\begin{array}[]{cc}17&-6i\\ 4i&8\\ \end{array}\right).$ (140)
By using (140), we have
$A_{1}-A_{1}^{*}=0,$
$A_{3}^{*}A_{2}-A_{2}^{*}A_{3}=0,$
$A_{3}-A_{3}^{*}=\begin{pmatrix}0&-2i\\ -2i&0\\ \end{pmatrix},$ (141)
$A_{1}^{*}A_{2}-A_{2}^{*}A_{1}=\begin{pmatrix}0&-2i\\ 2i&0\\ \end{pmatrix}.$ (142)
From (141) and (142), we see that Condition C (14) is not satisfied. On the other hand, the determinant of ${\bf f}_{3}$ is given by
$\det{\bf f}_{3}=z^{6}+8z^{5}+23z^{4}+61z^{3}+121z^{2}+71z+112.$
The approximate value of its roots are
$z_{1}=$ $-4.903,z_{2}=-2.961,z_{3}=-0.055-1.117i,$
$z_{4}=$ $-0.055+1.117i,z_{5}=-0.012-2.482i,z_{6}=-0.012+2.482i.$
Consequently, ${\bf f}_{3}$ is a Hurwitz matrix polynomial.
As indicated in [12], not every Hurwitz matrix polynomial is an HTM polynomial. In the following subsection, we propose a procedure to complete a given matrix polynomial into an HTM polynomial.
5.1 Completing a matrix polynomial
@@
was suggested as a polynomial that is a Hurwitz polynomial but not an HTM polynomial. The rational function $\frac{1}{z}\left(A_{1}z+A_{3}\right)\left(I_{2}z+A_{2}\right)^{-1}$ does not satisfy property (72).
The following result allows determining whether a polynomial $P_{n}$ that is not an HTM polynomial is in fact a Hurwitz polynomial by “completing” $P_{n}$ with another polynomial $Q_{n-1}$.
@@
Let $P_{n}(z)$ be $q\times q$ a monic polynomial of degree $n$. If there exists a $q\times q$ polynomial $Q_{n-1}(z)$ of degree $n-1$ such that
@@
From Theorem 3.2 f) and (75), there are polynomials $P_{1,n}$ y $Q_{1,n}$ such that $P_{n}(z)=(-1)^{n}P_{1,n}^{*}(-\bar{z})$ $Q_{n-1}(z)=(-1)^{n+1}Q_{1,n}^{*}(-\bar{z})$ and $P_{1,n}(z)$ is an orthogonal polynomial on $[0,+\infty)$ with respect to a certain positive measure $\sigma$ corresponding to the Stieltjes transform $-\frac{Q_{1,n}^{*}(\bar{z})}{P_{1,n}^{*}(\bar{z})}$. By [25, Theorem 1], the roots of the polynomial $\det P_{1,n}(z)$ lie in $[0,+\infty)$. Consequently, the roots of polynomial $\det P_{n}(z)=\det((-1)^{n}P_{1,n}^{*}(-\bar{z}))$ lie in $(-\infty,0]$. ∎
@@
Step 1. Let $P_{1,n}(z)=(-1)^{n}P_{n}^{*}(-\bar{z})$ (resp. $P_{1,n}(z)=(-1)^{n}P_{n}^{\intercal^{*}}(-\bar{z})$). By using (44), compute $Q_{1,n}(z)$ in terms of the coefficients of $P_{n}(z)$ (resp. $P_{n}^{\intercal}(z)$) and the moments or Markov parameters $s_{j}$, associated with some positive measure $\sigma$ on $[0,+\infty)$.
@@
We apply the completion procedure to the matrix polynomial (143). For $P_{3}^{\intercal}(z)$, we use the above algorithm together with Remark 5.12. With moments $s_{0}=\begin{pmatrix}2b&-b\\ -b&b\end{pmatrix}$, $s_{1}=\begin{pmatrix}4b&-b\\ -b&b\end{pmatrix}$ and $s_{2}=\begin{pmatrix}16b&-2b\\ -2b&2b\end{pmatrix}$, we calculate $Q_{1,3}(z)$; therefore, we have
$Q_{1,3}(z)=s_{0}z^{2}+z(s_{1}-A_{1}^{\intercal^{*}}s_{0})+s_{2}-A_{1}^{\intercal^{*}}s_{1}+A_{2}^{\intercal^{*}}s_{0}.$ (146)
From the Laurent expansion of $-\frac{P_{1,3}^{*}(\bar{z})}{Q_{1,3}^{*}(\bar{z})}$, we calculate the moments
@@
$B_{1}=\begin{pmatrix}2b&-b\\ -b&b\end{pmatrix},B_{2}=\begin{pmatrix}\frac{2732b}{109}&-8b\\ -8b&8b\\ \end{pmatrix},B_{3}=\begin{pmatrix}\frac{6826b}{109}&-11b\\ -11b&11b\end{pmatrix}.$ (148)
@@
In the scalar version, one can construct a sequence of scalar Hurwitz polynomials $(f_{j})_{j=1}^{\infty}$ using the sequence of scalar orthogonal polynomials on $[0,+\infty)$ $(p_{1,j})_{j=1}^{\infty}$ and $(p_{2,j})_{j=1}^{\infty}$ as in Definition 2.8; see [14].
@@
The proof of (149) and (150) readily follows from the interlacing properties of the roots of the orthogonal polynomials $p_{k,m}(z)$ and $\widetilde{p}_{k,m-1}(z)$ for $k=1,2$, and from the fact that $a_{0}a_{1}>0$, where $a_{0}$ and $a_{1}$ are the leading coefficients of the polynomial $f_{n}(z)=a_{0}z^{n}+a_{1}z^{n-1}+\cdots+a_{n}$; see the Hermite–Biehler theorem [5], [39], [58, page 46]. ∎

### v3

@@ 3 Bezoutian and Hurwitz-type matrix polynomials
$=$ ${\bf h}_{2m}^{*}(\bar{x}^{2})\left(x^{-2}I_{q},x^{-4}I_{q},\ldots\right)H^{(1)}_{\infty}\left(\bar{y}^{-2}I_{q},\bar{y}^{-4}I_{q},\ldots\right)^{*}{\bf h}_{2m}(y^{2}).$
@@
${\bf h}_{2m}^{(11)}(\xi)+{\bf h}_{2m}^{(12)}(\xi)=$ $\left(\bar{\xi}^{-2}I_{q},\bar{\xi}^{-4}I_{q},\ldots\right)^{*}{\bf h}_{2m}(\xi^{2}),$ (45)
${\bf h}_{2m}^{(21)}(\xi)+{\bf h}_{2m}^{(22)}(\xi)=$ $\left(\bar{\xi}^{-1}I_{q},\bar{\xi}^{-3}I_{q},\ldots\right)^{*}{\bf h}_{2m}(\xi^{2}),$ (46)
${\bf g}_{2m+1}^{(11)}(\xi)+{\bf g}_{2m+1}^{(12)}(\xi)=$ $\left(I_{q},\bar{\xi}^{-2}I_{q},\ldots\right)^{*}{\bf g}_{2m+1}(\xi^{2}),$ (47)
${\bf g}_{2m+1}^{(21)}(\xi)+{\bf g}_{2m+1}^{(22)}(\xi)=$ $\left(\bar{\xi}^{-1}I_{q},\bar{\xi}^{-3}I_{q},\ldots\right)^{*}{\bf g}_{2m+1}(\xi^{2}),$ (48)
@@
$\widehat{\bf h}_{2m}^{(11)},\,\widehat{\bf h}_{2m}^{(21)},\,\widehat{\bf g}_{2m+1}^{(11)}\quad\mbox{and}\quad\widehat{\bf g}_{2m+1}^{(11)}$
@@
$\widehat{\bf h}_{2m}^{(11)}(y)=\left(\bar{y}^{-2}{\bf h}_{2m}^{*}(\bar{y}^{2}),\bar{y}^{-4}{\bf h}_{2m}^{*}(\bar{y}^{2}),\ldots,\bar{y}^{-2m}{\bf h}_{2m}^{*}(\bar{y}^{2})\right)^{*}$ (54)
@@
${\bf h}_{2m}^{(11)}(y)=\left(\bar{y}^{-2}{\bf h}_{2m}^{*}(\bar{y}^{2}),\bar{y}^{-4}{\bf h}_{2m}^{*}(\bar{y}^{2}),\ldots,\bar{y}^{-2m}{\bf h}_{2m}^{*}(\bar{y}^{2}),0_{q},0_{q},\ldots\right)^{*}.$
@@
is called a matrix symmetrizer of $Q(x)$. In [42, page 455], the concept of symmetrizer is introduced for scalar polynomials. Additionally, let
@@
Let ${\mathscr{G}}_{2m}^{(k)}$, ${\mathscr{G}}_{2m+1}^{(k)}$ for $k=1,2$ be as defined in (37) and (38). Additionally, let ${\bf h}_{n}$, ${\bf g}_{n}$, $S$, $\widetilde{J}_{m}$, $H_{1,j}$, $H_{2,j}$, $F_{0}^{2k}(x)$ and $F_{1}^{2k-1}(x)$ be as in (5), (8), (55), (49), (13), (56) and (57), respectively.
@@
${\mathscr{G}}_{2m}^{(2)}(x,y)=$ $-F_{1}^{(2m-1)^{*}}(\bar{x})S({\bf h}_{2m}^{*})\widetilde{J}_{m-1}H_{1,m-1}\widetilde{J}_{m-1}S({\bf h}_{2m})F_{1}^{(2m-1)}(y).$ (59)
@@ 4 Hurwitness of Hurwitz-type matrix polynomials
In this section, using the explicit form of the Bezoutian matrix (71) and (72), we verify that an HTM polynomial ${\bf f}_{n}$ is indeed a Hurwitz matrix polynomial.
In [52], the Hurwitz property of HTM polynomials was studied. However, the proof of [52, Theorem 3.5] concerning the inertia of the matrix polynomial ${\bf f}_{n}$ for $n=2m$ is neither explicit nor complete, and a proof for the case $n=2m+1$ is not provided. See Remark 6.2.
Recall that the inertia of matrix polynomials, denoted by $\gamma({\bf f}_{n})$, provides the number of eigenvalues of ${\bf f}_{n}$ with positive, negative and real zero parts, respectively; see [43, page 410].
We rewrite Corollary 3.5 for the pair $(ix,iy)$ instead of $(x,y)$.
@@
$V_{[0,2m-1]}(x):=$ $\left(I_{q},\bar{x}I_{q},\bar{x}^{2}I_{q},\ldots,\bar{x}^{2m-2}I_{q},\bar{x}^{2m-1}I_{q}\right)^{\mathsf{*}},$
$V_{[0,2m]}(x):=$ $\left(I_{q},\bar{x}I_{q},\bar{x}^{2}I_{q},\ldots,\bar{x}^{2m-1}I_{q},\bar{x}^{2m}I_{q}\right)^{\mathsf{*}}.$
Furthermore, let ${\mathscr{T}}_{2m}$ (resp. ${\mathscr{T}}_{2m+1}$) be elementary block matrices such that
@@
$\begin{pmatrix}F_{0}^{(2m)}(x)\\ F_{1}^{(2m-1)}(x)\end{pmatrix}={\mathscr{J}}_{m-1}^{(m)}{\mathscr{T}}_{2m+1}V_{[0,2m]}(x).$ (76)
The following remark readily follows.
@@
Using (75) and (76), equalities (73) and (74) can be written as follows:
@@
Each HTM polynomial is a Hurwitz matrix polynomial.
@@
A $q\times q$ HTM polynomial ${\bf f}_{2}$ of degree $2$ is a Hurwitz matrix polynomial; it admits the representation
${\bf f}_{2}(z)=I_{q}z^{2}+s_{0}z+s_{0}^{-1}s_{1}.$
Remark 4.6.
[…]
For $n=2m+1$ (resp. $n=2m$), the points of discontinuity of the distribution function $\widetilde{\sigma}_{2m,\max}$ (resp. $\widetilde{\sigma}_{2m,\min}$) that correspond to the measure $\sigma_{2m,\max}$ (resp. $\sigma_{2m,\min}$) are given by the roots of the polynomial $\det(P_{1,m}^{*}(-\bar{z}))$ (resp. $\det(zP_{2,m}^{*}(-\bar{z}))$); see (27) and (28).
On the other hand, the polynomials (25) and (26), which are constructed using the moments $s_{j}$, can be obtained from an absolutely continuous distribution on $0,\infty)$ that corresponds to a positive measure on $0,\infty)$, as shown in the next example.
Example 4.7.
Let $\widetilde{\sigma}(x)=\begin{pmatrix}4-2e^{-x/2}&ie^{-x}\\ -ie^{-x}&2-e^{-x}\\ \end{pmatrix}$ be defined on $0,\infty)$. We construct the moments $s_{0}=\begin{pmatrix}2&-i\\ i&1\end{pmatrix}$, $s_{1}=\begin{pmatrix}4&-i\\ i&1\end{pmatrix}$, $s_{2}=\begin{pmatrix}16&-2i\\ 2i&2\end{pmatrix}$, $s_{3}=\begin{pmatrix}96&-6i\\ 6i&6\end{pmatrix}$, $s_{4}=\begin{pmatrix}768&-24i\\ 24i&24\end{pmatrix}$ of $\widetilde{\sigma}(x)$. The sequence $(s_{j})_{j=0}^{4}$ is a Stieltjes positive definite sequence. By using these moments and Definition 2.4, we derive $P_{2,2}$ and $Q_{2,2}$. From (24), we obtain the HMT polynomial
${\bf f}_{5}(z)=z^{5}I_{2}+A_{1}z^{4}+A_{2}z^{3}+A_{3}z^{2}+A_{4}z+A_{5},$ (82)
$A_{4}=\begin{pmatrix}\frac{1158}{37}&0\\ -\frac{936}{37}i&6\\ \end{pmatrix}$ and $A_{5}=\begin{pmatrix}\frac{292}{37}&-2i\\ 2i&2\\ \end{pmatrix}$. Therefore, by Theorem 4.3, the polynomial ${\bf f}_{5}$ is a Hurwitz polynomial.
Next, we paraphrase [5, Theorem 7.10].
Theorem 4.8.
Let $n$ be a natural number with $n\geq 2$, and let ${\bf f}_{n}$ be a matrix-valued polynomial. Let ${\bf h}_{n}$ (resp. ${\bf g}_{n}$) denote the ${\bf h}_{n}$-part (resp. ${\bf g}_{n}$-part) of ${\bf f}_{n}$. Furthermore, let $n=2m$ (resp. $n=2m+1$) with $m\geq 1$, and let $H_{1,m-1}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m-1}$) be constructed from the Hermitian Markov parameters of the polynomial ${\bf f}_{n}$. Then the matrix-valued polynomial ${\bf f}_{n}$ is an HTM polynomial if and only if the block Hankel matrices $H_{1,m-1}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m-1}$) are positive definite.
@@
a) the matrix $B_{x{\bf g}_{n}^{*},{\bf h}_{n}^{*}}(x{\bf g}_{n},{\bf h}_{n})$ is positive definite, and
b) the matrix $B_{{\bf g}_{n}^{*},{\bf h}_{n}^{*}}({\bf g}_{n},{\bf h}_{n})$ is negative definite.
@@
From Theorem 4.8, see that assertion a) and b) are valid. ∎
@@
In [42, Theorem 2], it is proven that a scalar polynomial $f_{n}$ is Hurwitz if and only if $B_{xg_{n}^{*},h_{n}^{*}}(xg_{n},h_{n})$ is negative definite and $B_{g_{n}^{*},h_{n}^{*}}(g_{n},h_{n})$ is positive definite. Note that the Markov parameters related to [42, Theorem 2] have the opposite sign compared with those associated with parts (a) and (b) of Proposition 4.9.
5 Completing a matrix polynomial to an HTM polynomial
As indicated in [9], not every Hurwitz matrix polynomial is an HTM polynomial. In this section, we propose a procedure to complete a given matrix polynomial into an HTM polynomial.
@@ 5 Completing a matrix polynomial to an HTM polynomial
was suggested as a polynomial that is Hurwitz but not an HTM polynomial. The rational matrix function $\frac{1}{z}\left(A_{1}z+A_{3}\right)\left(I_{2}z+A_{2}\right)^{-1}$ does not satisfy property (35). Moreover, as indicated in [9], the $2\times 2$ matrix coefficients $c_{j}$ of the expansion
$\frac{1}{z}\left(A_{1}z+A_{3}\right)\left(I_{2}z+A_{2}\right)^{-1}=-\frac{c_{0}}{z}-\frac{c_{1}}{z^{2}}-\frac{c_{2}}{z^{3}}$-$\cdots$
are not Hermitian matrices. Consequently, the polynomial $P_{3}$ cannot be investigated in the framework of HTM polynomials.
The following result allows one to determine whether a matrix polynomial $P_{n}$ that is not an HTM polynomial is in fact a Hurwitz polynomial by “completing,” mainly by perturbing $P_{n}$ with another polynomial $Q_{n-1}$. This is done by constructing a new HTM polynomial of degree $2n$.
@@
Let $P_{n}(z)$ be $q\times q$ a monic polynomial of degree $n$. If there exists a $q\times q$ polynomial $Q_{n-1}(z)$ of degree n1n-1 such that
@@
From (24) and (25), there are polynomials $P_{1,n}$ y $Q_{1,n}$ such that $P_{n}(z)=(-1)^{n}P_{1,n}^{*}(-\bar{z})$ $Q_{n-1}(z)=(-1)^{n+1}Q_{1,n}^{*}(-\bar{z})$ and $P_{1,n}(z)$ is an orthogonal polynomial on $0,+\infty)$ with respect to a certain positive measure $\sigma$ corresponding to the Stieltjes transform $\frac{Q_{1,n}^{*}(\bar{z})}{P_{1,n}^{*}(\bar{z})}$. By [22, Theorem 1], the roots of the polynomial $\det P_{1,n}(z)$ lie in $0,+\infty)$. Consequently, the roots of polynomial $\det P_{n}(z)=\det((-1)^{n}P_{1,n}^{*}(-\bar{z}))$ lie in $(-\infty,0]$. ∎
@@
Step 1. Let $P_{1,n}(z)=(-1)^{n}P_{n}^{*}(-\bar{z})$ (resp. $P_{1,n}(z)=(-1)^{n}P_{n}^{\intercal^{*}}(-\bar{z})$). By using (22), compute $Q_{1,n}(z)$ in terms of the coefficients of $P_{n}(z)$ (resp. $P_{n}^{\intercal}(z)$) and the moments or Markov parameters $s_{j}$, associated with some positive measure $\sigma$ on $0,+\infty)$.
@@
We apply the completion procedure to the matrix polynomial (83). Let the superscript $\mathsf{T}$ denote the transpose of a matrix. For $P_{3}^{\intercal}(z)$, we use the above algorithm together with Remark 5.2. With moments $s_{0}=\begin{pmatrix}2b&-b\\ -b&b\end{pmatrix}$, $s_{1}=\begin{pmatrix}4b&-b\\ -b&b\end{pmatrix}$ and $s_{2}=\begin{pmatrix}16b&-2b\\ -2b&2b\end{pmatrix}$, we calculate $Q_{1,3}(z)$; therefore, we have
$Q_{1,3}(z)=s_{0}z^{2}+z(s_{1}-A_{1}^{\intercal^{*}}s_{0})+s_{2}-A_{1}^{\intercal^{*}}s_{1}+A_{2}^{\intercal^{*}}s_{0}.$
From the Laurent expansion of $\frac{P_{1,3}^{*}(\bar{z})}{Q_{1,3}^{*}(\bar{z})}$, we calculate the moments
@@
$B_{1}=\begin{pmatrix}2b&-b\\ -b&b\end{pmatrix},B_{2}=\begin{pmatrix}\frac{2732b}{109}&-8b\\ -8b&8b\\ \end{pmatrix},B_{3}=\begin{pmatrix}\frac{6826b}{109}&-11b\\ -11b&11b\end{pmatrix}$.
@@
In the scalar version, one can construct a sequence of scalar Hurwitz polynomials $(f_{j})_{j=1}^{\infty}$ using the sequence of scalar orthogonal polynomials on $0,+\infty)$ $(p_{1,j})_{j=1}^{\infty}$ and $(p_{2,j})_{j=1}^{\infty}$ as in Definition 2.4; see [11].
@@
The proof of (87) and (88) readily follows from the interlacing properties of the roots of the orthogonal polynomials $p_{k,m}(z)$ and $\widetilde{p}_{k,m-1}(z)$ for $k=1,2$, and from the fact that $a_{0}a_{1}>0$, where $a_{0}$ and $a_{1}$ are the leading coefficients of the polynomial $f_{n}(z)=a_{0}z^{n}+a_{1}z^{n-1}+\cdots+a_{n}$; see the Hermite–Biehler theorem [3], [33], [48, page 46]. ∎